Calculate current through individual resistors in the following circuit diagram50 Ω 20 Ω 30 Ω + 12V 12 13 15 Ω 20 Ω 30 V (14) i4 15 - 80 Ω 20 V 60 Ω
Answer #1

in 50 M I (1,-12) +30V ¥12-13) dir 202 30 12V Dit I wi 13 니 15 I go ao Dit lis 80 20V 60 25) LOOP (LD KuL in 30 = 150i, -2012from ® 5 숴 is - 20 +2013 80 is=1 / 13 - 12 = 5013-2012–2014** -12=5083-3012 +5-513 17 4583-30i2 from equations 4 --> 30= 15Resistor cuortfA) 3 Or 14 Amy sor 202 o. 349 Amp 302 0.308 amps -0.149 amy 80 o.1991A Gor -0.379ame 20(120) -0.138 A * CS Sca